OpenStudy (anonymous):

Rory uploaded 1/5 of the files in his folder to his website before lunch, and another 3/7 after lunch. How much more of the files were uploaded after lunch than before lunch ?

OpenStudy (anonymous):

3/7 - 1/5 = 15/35 - 7/35 = 8/35 So it's 8/35 of the files

OpenStudy (anonymous):

Thank you so much . I am helping my 5th grader with his homework , and clearly I am not Smarter Than A Fifth Grader!!!